• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Holmstrom

    Log entry time 08:03:04 on July19,2005

    Entry number 148577

    keyword=Owl Shift Summary

    00:00 Start of shift, work is ongoing in the injector - no beam.
    00:30 MCC called no estiamte.
    03:45 Lisa went over to MCC and they think it might be an hour more.
    05:00 MCC called we will get a burp of beam and then they will work on the Hall 
    C line for about 40-45 minutes. 
    06:20 we start getting some beam:
    setpoint   spot++   actual
    3.8x4.75   2x2       3x3
    7.5x8.7    3.5x3
    7.5x10.0   3.5x3
    7.5x11.0   3.5x3.5   5x5
    11.3x14.0  4.5x4.5
    07:10 We switch to the He4 target.
    07:20 The ion chamebrs trip with the largest raster at 7uA.
    07:25 MCC having trouble.
    08:00 Shift ends